|-2-x|is greater than or equal to 4
Answer by Nate(3500)

|-2-x|is greater than or equal to 4
|-2-x| >= 4
-2-x >= 4
-x >= 6
x =< -6
or
-2-x =< -4
-x =< -2
x >=2